The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 42.7 for a sample of size 519 and standard deviation 19.3.

Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 95% confidence level).

Enter your answer as a tri-linear inequality accurate to one decimal place (because the sample statistics are reported accurate to one decimal place).

< μ<
Answer #1

Date: 1 Solution - Given that, x = 42.7 6 = 19:3 n=519 At as % confidence level, L=1-95 % =1-0-95 = 0.05 -0.05 0-025 Faiza 20